AutoFido Manufacturing Ltd. is a well-known manufacturer of high tech products. The president of AutoFido has decided he would like to manufacture an AI droid to take care of pets. It is rather rudimentary and will be able to walk, feed and entertain man's best friend. Based on his experience in the industry, he believes the new product will be a popular item for four years, after which it will be obsolete due to more advanced technology.

One time R & D costs are expected to be $2,400,000. Fixed production costs are anticipated to be $12,500 per month. Fixed marketing costs will be $20,000 per month for the first three years, dropping to $5,000 per month for the last year. Fixed distribution costs will be $20,000 per month for the first three years, and $10,000 per month for the last year.

Variable production costs will be $3,500 per unit for all four years. Variable distribution costs will be $50 per unit for the first three years, and $25 per unit for the last year. There are no variable marketing costs.

AutoFido expects to sell 400 units per month for the first three years, dropping to 100 units per month in the last year before the product becomes obsolete. The anticipated selling price is $5,000 per unit.

Required:

a.         Prepare a projected life cycle income statement for the new product line.

b.         How much profit, on average, will AutoFido make on each Falcon? (Round your answer to the nearest cent, i.e. 2 decimal places)

c.         If research shows the highest price customers will be willing to pay is $3,999 per unit, should AutoFido proceed with the AI droid, and why?

d.         If the only cost AutoFido can reduce is Variable Production costs, and they require full cost + 10%, what must be the new variable production cost per unit?
Answer #1

Answer :-

(a) Projected Life Cycle Income Statement for the New Product Line

Particulars Year 1 Year 2 Year 3 Year 4 Total
Sales (W.N. 1) 24000000 24000000 24000000 6000000 78000000
(-) R & D Costs (W.N. 2) 600000 600000 600000 600000 2400000
(-) Fixed Production Costs (W.N. 3) 150000 150000 150000 150000 600000
(-) Fixed Marketing Costs (W.N. 4) 240000 240000 240000 60000 780000
(-) Fixed Distribution Costs(W.N. 5) 240000 240000 240000 120000 840000
(-) Variable Production Costs (W.N. 6) 16800000 16800000 16800000 4200000 54600000
(-) Variable Distribution Costs (W.N. 7) 240000 240000 240000 30000 750000
= Anticipated Profit 5730000 5730000 5730000 840000 18030000

Working Notes:-

1.) Sales for first three years = 400*12*5000 = $ 24000000 & Sales for the fourth year = 100*12*5000 = $ 6000000

2.) R & D Costs total $ 2400000, so per year = 2400000/4 = $ 600000

3.) Fixed Production Costs $ 12500 per month, so 12500*12 = $ 150000 per year

4.) Fixed Marketing Costs for first three years = 20000*12 = $ 240000 & for fourth year = 5000*12 = $ 60000

5.) Fixed Distribution Costs for first three years = 20000*12 = $ 240000 & for fourth year = 10000*12 = $ 120000

6.) Variable Production Costs $ 3500 per unit, so for first three years = 3500*400*12 = $ 16800000 and for fourth year = 3500*100*12 = $ 4200000

7.) Variable Distribution Costs for first three years = 50*400*12 = $ 240000 and for fourth year = 25*100*12 = $ 30000

(b) Average Profit made on each falcon :-

Total number of units during the four years = (400*12*3) + (100*12*1) = 14400+1200 = 15600

Projected profit = $ 18030000

Average Profit = 18030000/15600 = $ 1155.77

(d) Profit required = Full Cost + 10%

So, sales price of $ 3999 includes cost of $ 3635.45 (3999/110*100) which is total of fixed & variable cost per unit.

now, fixed cost per unit = 4620000/15600 = $ 296.15

Hence variable cost required = 3635.45-296.15 = $ 3339.3, in which variable distribution costs = $ 48.08 Hence variable production cost per unit = $ 3291.22

Current variable production cost per unit = 54600000/15600 = $ 3500

Reduction required in variable cost per unit = 3500 - 3291.22 = $ 208.78

Hence new variable production cost per unit = $ 3291.22